# A lensed FRB candidate in the first CHIME/FRB Catalogue and its potential implications

### v1

@@ 2.1 Method
where $\bar{I}$ represents the mean value of $I_{t}$. According to Ji et al. (2018), a lensed signal will exhibit three spikes at $\delta t=-\Delta t,0,\Delta t$ with amplitude ratios $R/(R^{2}+1):1:R/(R^{2}+1)$, where $\Delta t$ and $R$ represent the time delay and the flux ratio between the lensed peaks, respectively.
For real data, the ACF typically shows a spiky structure, meaning $\delta t=-\Delta t,0,\Delta t$ are not the only peaks. Therefore, we need to define a threshold to distinguish lensing-induced spikes from noise. Here we apply Gaussian smoothing to the ACF, choosing $1/3$ as the standard deviation of the one-dimensional Gaussian kernel to minimize the influence of adjacent points on the central point. The dispersion between the ACF and the smoothed function $F(k)$, following Wang et al. (2021), is defined as:
@@ 2.2 Data
The data used in this study is from the first CHIME/FRB catalogue (Masui & Chime/Frb Collaboration, 2021), which contains 536 FRBs. The parameters produced in this catalogue indicate that 42 of these FRBs exhibit multi-peak structures, suggesting that some could be potential candidates for lensed bursts. Due to the existence of noise in the raw data, the result from the autocorrelation method could be significantly impacted. Therefore, we also applied this method to the model-fitted data provided in the catalogue.
@@ 2.4 Simulation
For signals that have passed the initial screening criteria, the existing conditions may still not be robust enough to confirm lensing. To strengthen our results, we also employ signal simulation. By controlling the parameters of the substructure to be consistent and simulating results that are most similar to the observed signals, we can determine whether these signals could plausibly originate from lensed systems.
@@ 3.1 Model introduction
$\Delta t=\frac{4GM}{c^{3}}\left(1+z_{l}\right)\left[\frac{y_{2}s}{2\left(1-\gamma\right)}+\ln\left(\frac{s+y_{2}}{s-y_{2}}\right)\right]$ (4)
@@
$R=\frac{\left(y_{2}^{2}+2+y_{2}s\right)\left(y_{2}s-4\gamma\right)+8\gamma^{2}+2\gamma y_{2}^{2}}{\left(y_{2}^{2}+2-y_{2}s\right)\left(y_{2}s+4\gamma\right)-8\gamma^{2}-2\gamma y_{2}^{2}},$ (5)
@@
$s=\sqrt{y_{2}^{2}+4\left(1-\gamma\right)}.$ (6)
@@
$\Delta t=\frac{4GM}{c^{3}}\left(1+z_{l}\right)\left[\frac{y_{1}s}{2\left(1+\gamma\right)}+\ln\left(\frac{s+y_{1}}{s-y_{1}}\right)\right]$ (7)
@@
$R=\frac{\left(y_{1}^{2}+2+y_{1}s\right)\left(y_{1}s+4\gamma\right)+8\gamma^{2}-2\gamma y_{1}^{2}}{\left(y_{1}^{2}+2-y_{1}s\right)\left(y_{1}s-4\gamma\right)-8\gamma^{2}+2\gamma y_{1}^{2}},$ (8)
@@
$s=\sqrt{y_{1}^{2}+4\left(1+\gamma\right)}.$ (9)
@@ 3.2 Constraints on lens mass
As discussed in Section 3, for FRB 20190308C, we can derive constraints on the lens mass using its flux ratio $R=2.00$ and time delay $\Delta t=8.85\,\rm ms$, as shown in Figure 4. The blue and orange lines in the figure are obtained from Equations (7) and (8), the red line from Equation (10), and the green line from Equations (4) and (5). The black dashed line represents the lens mass constraints for $\gamma=0.01$. For $R=2.00$, $\Delta t=8.85\,\rm ms$ and $\gamma=0.01$, the lens mass limit obtained through this method is $\left[560M_{\odot},823M_{\odot}\right]$ and $\left[2980M_{\odot},11280M_{\odot}\right]$.
Figure 4: The range of lens mass constraints corresponding to the variation of $\gamma$ from 0 to 0.1. The blue and orange curves are given by Equations (7) and (8), the red curve is obtained by Equation (10), and the green curve can be determined using Equations (4) and (5). The black vertical dashed line represents the typical condition when $\gamma=0.01$ and the constrained lens mass range is $\left[560M_{\odot},823M_{\odot}\right]$ and $\left[2980M_{\odot},11280M_{\odot}\right]$.
@@ 4 Discussion and Conclusions
If a lensed signal shows a pulse profile similar to FRB 20190308C, where the first peak is weaker than the second one, this cannot be explained by the point-mass lens model. Therefore, we used the Chang–Refsdal lens model and obtained a range of lens mass as shown in Figure 4. The mass range varies with different values of $\gamma$. For a typical of $\gamma=0.01$, the lens mass is constrained to be $\left[560M_{\odot},823M_{\odot}\right]$ and $\left[2980M_{\odot},11280M_{\odot}\right]$.
According to Liao et al. (2022), to identify lensed FRB signals, it is necessary to find at least two pairs of bursts with the same time delay, flux ratio and DM difference, or to observe inconsistent source positions. Therefore, future observations with longer durations and higher precision will be required to find convincing lensed signals. What we present in this article are only potential candidates for lensed FRB signals.

### v2

@@ 2.1 Method
where $\bar{I}$ represents the mean value of $I_{t}$. According to Ji et al. (2018), a lensed signal will exhibit a spike at $\delta t=\Delta t$, where $\Delta t$ represent the time delay between the lensed peaks. For real data, the ACF typically shows a spiky structure, meaning $\delta t=\Delta t$ may not be the only peak. Therefore, we need to define a threshold to distinguish lensing-induced spikes from noise. Here we apply Gaussian smoothing to the ACF, choosing $1/3dt$ as the standard deviation of the one-dimensional Gaussian kernel to minimize the influence of adjacent points on the central point, where $dt=0.98\,\rm ms$ is the time resolution of the data. The dispersion between the ACF and the smoothed function $F(k)$, following Wang et al. (2021), is defined as:
@@ 2.2 Data
The data used in this study is from the first CHIME/FRB catalogue (Masui & Chime/Frb Collaboration, 2021), which contains 536 FRBs. The parameters produced in this catalogue indicate that 42 of these FRBs exhibit multi-peak structures, suggesting that some could be potential candidates for lensed bursts. Due to the existence of noise in the raw data, the result from the autocorrelation method could be significantly impacted. Therefore, we also applied this method to the model-fitted data provided in the catalogue. As long as either the original data or the model-fitted data of a signal passes the screening, we would add it to the list of potential candidates.
@@
To address the possibility that model-fitted data might lose fine temporal structure present in the original data, we calculated the residuals between the model-fitted and the original data. To investigate the impact of this on the search, we used the same method to calculate the autocorrelation function of the residuals, to search for temporal structures that might have not been captured by the model fit. The results indicate that there is no significant correlation in the residual autocorrelation for these 42 signals. If the background noise is sufficiently clean, correlated fine structures might still be present in the residuals after model smoothing of a lensed signal. Therefore, a null search in the residual auto-correlation might indicate a non-detection of lensing. However, due to the complex nature of background noise components, analysing the signal solely through its dynamic spectrum is challenging. Hence, this paper focuses specifically on the lensing possibility of the model-fitted pulsing signals.
@@ 2.4 Simulation
For signals that have passed the initial screening criteria, the existing conditions may still not be robust enough to confirm the possibility of lensing.
According to Masui & Chime/Frb Collaboration (2021), the model-fitted data is treated as a superposition of different components. A signal consisting of $N$ components can be represented as follows:
$S=\sum_{i=0}^{N}A_{i}\times F_{i}\times T_{i},$ (3)
where $A_{i}$ is the overall amplitude of the $i$th burst component; $F_{i}\left(\gamma_{i},r_{i}\right)$ is the time-independent spectral energy distribution and $T_{i}\left(DM,t_{arr,i},w_{i},\tau\right)$ is the temporal shape of the burst. Although we attempted to identify similar components in the time series structure by calculating the autocorrelation function, there are slight differences in various parameters of these similar components, such as signal width ($w_{i}$) and spectral indices ($\gamma_{i}$ and $r_{i}$) of the spectral energy distribution since these parameters were uniquely defined for each of the $N$ components.
A lensed system theoretically requires these parameters to remain consistent across different substructures, while deviations always exist in practice. These differences may be inherent to the signal components themselves, or they may be caused by noise interference or fitting errors. We cannot distinguish how these differences arise, but we can simulate lensed FRB signals by controlling these parameters to remain consistent within the substructures of the signal, in order to investigate whether lensing effects can produce similar dynamic spectra of the candidates.
To strengthen our results, we also employ signal simulation. By controlling the parameters of the substructure to be consistent and simulating results that are most similar to the observed signals, we can determine whether these signals could plausibly originate from lensed systems.
@@ 3.1 Model introduction
$\Delta t=\frac{4GM}{c^{3}}\left(1+z_{l}\right)\left[\frac{y_{2}s_{2}}{2\left(1-\gamma\right)}+\ln\left(\frac{s_{2}+y_{2}}{s_{2}-y_{2}}\right)\right]$ (5)
@@
$R=\frac{\left(y_{2}^{2}+2+y_{2}s_{2}\right)\left(y_{2}s_{2}-4\gamma\right)+8\gamma^{2}+2\gamma y_{2}^{2}}{\left(y_{2}^{2}+2-y_{2}s_{2}\right)\left(y_{2}s_{2}+4\gamma\right)-8\gamma^{2}-2\gamma y_{2}^{2}},$ (6)
@@
$s_{2}=\sqrt{y_{2}^{2}+4\left(1-\gamma\right)}.$ (7)
@@
$\Delta t=\frac{4GM}{c^{3}}\left(1+z_{l}\right)\left[\frac{y_{1}s_{1}}{2\left(1+\gamma\right)}+\ln\left(\frac{s_{1}+y_{1}}{s_{1}-y_{1}}\right)\right]$ (8)
@@
$R=\frac{\left(y_{1}^{2}+2+y_{1}s_{1}\right)\left(y_{1}s_{1}+4\gamma\right)+8\gamma^{2}-2\gamma y_{1}^{2}}{\left(y_{1}^{2}+2-y_{1}s_{1}\right)\left(y_{1}s_{1}-4\gamma\right)-8\gamma^{2}+2\gamma y_{1}^{2}},$ (9)
@@
$s_{1}=\sqrt{y_{1}^{2}+4\left(1+\gamma\right)}.$ (10)
@@ 3.2 Constraints on lens mass
As discussed in Section 3, for FRB 20190308C, we can derive constraints on the lens mass using its flux ratio $R=0.5$ and time delay $\Delta t=8.85\,\rm ms$, as shown in Figure 4 based on the chosen value of $\gamma$. Here we only consider the condition when $10<\gamma<$ 1 following the results presented in Chen et al. (2021a). The blue line in the figure are obtained from Equations (5) and (6), and the orange line from Equation (11). It is clear that the range of the combined mass $M\left(1+z_{l}\right)$ varies with different values of $\gamma$. The black dashed line taken as an example represents the lens mass constraints for $\gamma=0.01$. For $R=0.5$, $\Delta t=8.85\,\rm ms$ and $\gamma=0.01$, the limit of the combined lens mass is $\left[4277M_{\odot},11280M_{\odot}\right]$.
Figure 4: The range of lens mass constraints corresponding to the variation of $\gamma$ from 0 to 0.1. The blue is given by Equations (5) and (6), and the orange line is obtained by Equation (11). The black vertical dashed line represents the condition when $\gamma=0.01$ and the combined lens mass range is $\left[4277M_{\odot},11280M_{\odot}\right]$
@@ 4 Discussion and Conclusions
If a lensed signal shows a pulse profile similar to FRB 20190308C, where the first peak is weaker than the second one, this cannot be explained by the point-mass lens model. Therefore, we used the Chang–Refsdal lens model and obtained a range of the combined lens mass as shown in Figure 4. The mass range varies with different values of $\gamma$. If $10<\gamma<$ 1, for a chosen value of $\gamma=0.01$, the combined lens mass can be constrained to be $\left[4277M_{\odot},11280M_{\odot}\right]$. The case when $\gamma>1$ is more complex and will be discussed in future work.
It is worth noting that the possibility of FRBs generating similar structures both temporally and spectrally through random processes always exists. Such coincidentally generated similar structures and those produced by gravitational lensing are almost indistinguishable based solely on the dynamic spectrum of the signal. Therefore, this paper seeks only potential lensed FRB candidates rather than definitive lensed FRBs. According to Liao et al. (2022), to identify lensed FRB signals, it is necessary to find at least two pairs of bursts with the same time delay, flux ratio and DM difference, or to observe inconsistent source positions. Therefore, future observations with longer durations and higher precision will be required to find convincing lensed signals. What we present in this article are only potential candidates for lensed FRB signals.